OpenStudy (anonymous):

Show that tan15 degrees = 2 - sqrt(3) by using the double angle formula.

OpenStudy (anonymous):

\[\tan 15=2-\sqrt{3}\] by using double angles

OpenStudy (whpalmer4):

I think that's the wrong formula — you want the half angle formula. \[\tan^2 u = \frac{1-\cos(2u)}{1+\cos(2u)}\] \[\tan u = \sqrt{ \frac{1-\cos(2u)}{1+\cos(2u)}}\]You know the value of \(\cos(2u)\) if \(u=15\)...

OpenStudy (anonymous):

\[\tan15= \tan(45-30)\] expand that

OpenStudy (anonymous):